OpenStudy (anonymous):

Which of the following is the correct graph of the solution to the inequality −8 greater than or equal to −5x + 2 > −38? number line with an open dot at 2 with shading to the left and another open dot at 8 with shading to the right. number line with an open dot at 8 and at 2 with shading in between. number line with a closed dot on 2 and an open dot on 8 and shading in between. number line with closed dots at 2 and 8 with shading between the 2 and the 8.

HanAkoSolo (jamierox4ev3r):

\[-8\ge-5x+2>-38\]

HanAkoSolo (jamierox4ev3r):

is this what you are asking?

OpenStudy (anonymous):

the second one

HanAkoSolo (jamierox4ev3r):

wait what?

OpenStudy (anonymous):

which one is the correct graph of the solution

OpenStudy (texaschic101):

-8 >= -5x + 2 -8 - 2 >= -5x -10/-5 <= x (flip the sigh when dividing by a negative) 2 <= x -5x + 2 > -38 -5x > -38 - 2 -5x > -40 x < -40/-5 x < 8 solution is : closed dot on 2, open dot on 8, shading in between